OpenStudy (anonymous):

The difference between the digits in a two-digit numeral is 3. The units' digit is twice the tens' digit. Find the numeral.

OpenStudy (tkhunny):

Half the battle is in the DEFINITION. We have two digits. Only 0, 1, 2, ..., 9 are acceptable results. T = the tens digit U = the units digit We're almost done. "The difference between the digits in a two-digit numeral is 3." Actually, this is a very tricky hint. We don't know which digit is the greater of the two. Let's go read the other hint and see if it tells us. "The units' digit is twice the tens' digit." Ah, then U = 2*T and U is greater than T. Now, we can translate the difference hint. "The difference between the digits in a two-digit numeral is 3." U - T = 3 Okay, we have U = 2T and U - T = 3. How shall we find values for U and T?

OpenStudy (anonymous):

I need help on finding the solution to this problem. I only could partially understand it.